# Scalable Ab Initio Electronic Structure Methods with Near Chemical Accuracy for Main Group Chemistry

### v1

@@ 3.3 CCSD(T) and DLPNO-CCSD(T)
3.4 Basis Sets
We use the following basis sets: aug-cc-pVXZ-DK (X = T, Q) for atoms with atomic number less than or equal to that of oxygen, and aug-cc-pCVXZ-DK (X = T, Q) for fluorine and heavier, obtained using the Basis Set Exchange database.54, 55, 56, 57, 58, 59, 60 This choice is motivated by the documented improvement of basis set convergence when using core-valence or tight-d functions in the basis set for second-row elements (Na-Cl)61, 62, 63, 64 as well as fluorine65. While the split-valence aug-cc-pVXZ basis sets are not designed for core-valence correlation 66, we find that all-electron calculations using these same basis sets can reach a respectable (especially for CCSD(T)) albeit overall inferior accuracy (especailly for AFQMC) to the frozen core calculations supplemented with MP2 core corrections, as discussed briefly in Section IV, and in more in detail in SI Section 5.
We extrapolate all single point energies to the complete basis set (CBS) limit according to the method of Neese and Valeev67 for T/Q extrapolation for both Hartree-Fock and correlation energy, with $\alpha$ and $\beta$ matching the basis set used. We use the same coefficients $\alpha$ and $\beta$ for the core-valence sets aug-cc-pCVXZ-DK as the corresponding aug-cc-pVXZ-DK basis sets, where we use the aug-cc-pVXZ coefficients reported by Neese and Valeev. This CBS procedure is used by all the methods investigated, with the exception of the more expensive CCSD(T) and W-AFQMC where we use alternative schemes (see SI Section 3 and Section 6).
3.5 Frozen Core Corrections
@@
3.6 Relativistic Effects
Scalar relativistic effects are included through the DKH268, 69, 70 Hamiltonian for DLPNO-CCSD(T) and CCSD(T), and x2c71 for ph-AFQMC, as x2c is not implemented in ORCA and DKH2 is not implemented in PySCF. The MP2 core-valence corrections follow the same relativistic corrections for each method, respectively (MP2 is carried out in ORCA for correcting DLPNO-CCSD(T) and CCSD(T) corrections and in PySCF for ph-AFQMC corrections). Spin-orbit corrections to atomic energies are applied using the values from Curtiss et al. 1.
3.7 Heat of Formation and Atomization
Atomization energies ($\sum E_{\text{atoms,{g}}}-E_{\text{molecule,{g}}}$) and heats of formation are calculated according to the method in Ref.1, 72, 73, including geometry optimization and thermochemical properties (ZPE, enthalpy, internal energy) using the DFT functional B3LYP 74, 75 and basis set 6-31G* 76, 77, 78, 79 using the Jaguar software package 80 with the maximum available grid point density. A few molecules required a higher-level geometry optimization (see SI Section 7). After optimization, DFT single-point energies were calculated with Jaguar. We note that the first step towards obtaining the heat of formation at 298 K is the atomization at 0 K, and the molecule temperature corrections (from DFT), atom temperature corrections (from experiment) and energy to change the atomic states from gas to standard state (from experiment, i.e. heat of formation of the single atom in the gaseous state) are added to achieve the heat of formation. Experimental values for the atomic heats of formation and temperature corrections are the same as that used by Curtiss et al.1, with the exception of the sulfur atom heat of formation where we use 66.18 kcal/mol from ATcT 40 instead of 65.66 kcal/mol.
@@ 3.7 Heat of Formation and Atomization
and with the addition of aformentioned temperature correction terms from DFT ($H^{298\text{ K}}-H^{0\text{ K}}$ for the molecule) and experiment ($H^{298\text{ K}}-H^{0\text{ K}}$ for the atom), we obtain $\Delta_{f}H(298\,\text{K})$. Although these corrections are approximate (even though we expect the errors to be small), they cancel out when obtaining the deviation from the converted W4 heat of formation reference value as the same corrections are applied to the calculated atomization energy of the molecule. Effectively, the deviation $D$ becomes
@@
3.8 Atomic Energies
The treatment of atoms is essential in achieving the targeted accuracy. An explicit near-exact treatment of core-valence correlation on par with valence correlation, as is done in W4-17, requires expensive core-valence corrections. In AFQMC, similar to other projection QMC approaches, the description of core-valence correlations requires onerous convergence of time-step errors. Furthermore, it has been shown that AFQMC atom energies can be difficult to calculate 35. Hence, we converge the atomic energies with a large number of determinants in W-AFQMC. A less expensive and simpler alternative is to fit the values of the atoms to the experimental data, which benefits from cancellation of errors. We use this approach for the other benchmark methods. For small molecules, a relatively inexpensive version of the first approach can be shown to work quite well, however, for larger molecules, any difference in molecule versus atom accuracy for a given method is compounded.
@@ 3.8 Atomic Energies
where $\Delta_{f}H_{\text{expt},\text{molecule}}$ and $\Delta_{f}H_{\text{calc}}$ are the experimental and calculated heats of formation for that molecule, and $N_{\text{atom}}$ is the number of the atom constituting the molecule. The energy of the atom, $E_{\text{atom}}$, is the fitting variable, and $c_{\text{molecule}}$ are the constant terms (for each molecule, i.e. independent of $E_{\text{atom}}$) such as the temperature terms, energy of the molecule, and heat of formation of the atoms, used when we minimize the left-hand side of the equation above. We have 189 such correlated equations for the 189 molecules in the G2 and G3 sets.
The initial guess for atom energies in the atomic energy fit is the ab initio atom energies obtained through each method respectively. It is worth noting that a small error (i.e. slight imperfection in cancellation of error between atom and molecule energy) in the ab initio atom energies is multiplied by the number of atoms in the molecule. While we do add spin-orbit corrections to atomic energies using the values from Curtiss et al. 1, this only applies to the initial guess and such atom-related corrections will be encompassed in the final fitted atom energy.
@@
4 IV. Results and Discussion
4.1 Overall Performance of CCSD(T) and AFQMC-based Methods
We begin by evaluating the overall performance of the four methods discussed in Section III: CCSD(T), DLPNO-CCSD(T), AFQMC I, and AFQMC II, across our chosen datasets. Figure 1 displays the root mean square deviation (RMSD) in the enthalpies of formation calculated using these methods, as detailed in Section III. For the combined dataset (represented by the far-right bar in Figure  1), the RMSD values for all methods fall within 1 kcal/mol, and as we will see below, all four methods have a very small number of outliers with deviations from benchmark experiments or computations greater than 2 kcal/mol. We conclude that for main group chemistry, the localized (and hence scalable) versions of both coupled cluster and AFQMC achieve our target of reliably obtaining near-chemical accuracy for chemical transformations, sufficient for elucidating chemical reaction mechanisms in complex systems. Results along these lines have been presented previously for DLPNO-CCSD(T) (although not for as large and extensively curated a data set involving experimental and high-level theoretical references), but not for AFQMC. The present exercise establishes AFQMC as a robust alternative benchmark quantum chemical methodology, albeit at a higher computational cost than DLPNO-CCSD(T). See SI Section 9 for examples of the computational costs of each method.
@@ 4.1 Overall Performance of CCSD(T) and AFQMC-based Methods
Secondly, the most difficult data set for all methods is, unsurprisingly, but not guaranteed, the subset of 10 cases that we have identified as “multireference” (MR). We classify molecules as multireference based on a set of diagnostics developed by Karton et. al.3, 4, as discussed in more detail in Section 10 of the Supplementary Material. For the coupled-cluster based methods, only one molecule, ozone, stands out as displaying an error in excess of 2 kcal/mol. Despite formally being a single reference methodology, the treatment of electron correlation via CCSD(T) appears to be powerful enough to handle many wavefunctions with nontrivial multireference character. DLPNO-CCSD(T) here displays a noticeable (although not large) degradation from full CCSD(T). For AFQMC, an improvement is obtained in the treatment of MR molecules by upgrading the trial function in the AFQMC II approach. Details of results for each quantum chemical method for all of the MR cases can be found in the SI Section 10.
@@
We next analyze the outliers observed across the different data sets and methods, the number of which are summarized in Figure 2 below. A number of interesting points can be made regarding the outliers, which are enumerated along with results for the various computational methods in Table 2. Firstly, neither AFQMC I, AFQMC II nor the two coupled cluster based methods have any outliers for the W4-11 and W4-17 data sets, where the reference values are taken from ultrahigh level computation. Secondly, while CCSD(T) and DLPNO-CCSD(T) have a larger number of outliers than AFQMC II based on our (somewhat arbitrary) cutoff for the deviation from the reference of 2 kcal/mol, a perusal of the data in Table 2 shows that, with the exception of ozone (where we believe that the multireference character is great enough to cause significant errors in CCSD(T) and DLPNO-CCSD(T), which can then be reduced via a large trial function in AFQMC II), the computational results for the five remaining molecules are closer to each other than they are to the experimental reference data. A likely interpretation of the results is that the experiments have a residual error of a few kcal/mol (possibly as large as 3-4 kcal/mol for 3-butyn-2-one), and that in fact, the reliability of our scalable benchmark methods is higher than what is suggested in Figure 2. In contrast to a number of other cases that initially appeared to be outliers but were resolved by newer experiments, as discussed in Section II, we were unable to find any relatively recent experimental data. Note that the value of having two distinct computational methods which can be compared, suggested in the Introduction, is already manifested in this analysis. Without the complementary AFQMC I and AFQMC II results, one might conclude that CCSD(T) has occasional outliers even for single-reference main group molecules and that higher-order treatments are required to achieve even the 2 kcal/mol accuracy threshold that we have set.
@@
4.2 Comparison of CCSD(T) and DLPNO-CCSD(T) Results
@@
4.3 Detailed Discussion of ph-AFQMC Methodology
In contrast to our coupled cluster based calculations, for which we utilize well established methods implemented by the Neese group in ORCA, the scalable AFQMC I protocol discussed above required significant novel methodology development. We therefore discuss AFQMC I development and implementation in detail in what follows. Specification of the AFQMC II protocol given AFQMC I as a starting point is straightforward, using the general procedure described in Section III.
@@ 4.3 Detailed Discussion of ph-AFQMC Methodology
Table 4: The top outliers from AFQMC 0 protocol. Deviations (($\Delta_{f}H_{\text{expt}}(298\,\text{K})-\Delta_{f}H_{\text{calc}}(298\,\text{K})$) for G2/G3 and $\Delta_{a}E_{\text{W4}}(0\,\text{K})-\Delta_{a}E_{\text{calc}}(0\,\text{K})$ for W4) are listed in kcal/mol with statistical errors in parentheses. After the first CI is performed with an active space (AS) based on orbital maps to the atoms of the molecules (refer Table S4) that returns the ‘First CI AS’ listed, the second AS (shown here as ‘TZ final AS’ and ‘QZ final AS’, as the NOONs have a slight basis set dependency due to approximations such as the SHCI solver) is chosen from those orbitals from the first AS that have NOONs of between 0.01 and 1.99. The final number of determinants is set by the number of determinants required to get to 99.5% saturation of the CI coefficients.
@@
Under circumstances where the reference is unknown, typically having a converged energy with respect to, for example, determinants30 (See SI Section 3 for an example) gives confidence in the AFQMC benchmark value unless the CI expansion is qualitatively wrong. This process can become expensive, and based on our heuristic we have observed some guidelines for which type of trial and whether convergence is necessary for certain types of molecules. Observing the AFQMC 0 outliers, a few categories of molecules stand out: i) multireference molecules ii) small (< 4 atoms) molecules containing Li, F, or S atoms, iii) conjugated systems, iv) strained systems, and v) halogen oxoacids. The only exceptions to these are diethyl ketone (deviation 2.22 kcal/mol) and HNCO (deviation 2.63 kcal/mol). After an expanded valence space to two instead of one shell, and a 0.001 instead of 0.01 NOON threshold, the remaining real outliers as discussed above mostly fall into the multireference category. Therefore, for main group thermochemistry, for the above categories of outliers (except multireference), we recommend AFQMC 0 with these alternative thresholds. On the other hand, we still recommend that multireference molecules be converged with respect to the active space size and number of determinants.
@@
The G2 and G3 test sets and the W4 sets are chosen on account of the readily available and accurate reference values. Further critical investigation of scalable benchmark methods such as DLPNO-CCSD(T) and AFQMC for larger systems is valuable, with the difficulty of such investigation being the lack of accurate experimental references and the computational expense of generating benchmark-level calculations for large datasets of such systems. Regardless, the current studies in the literature are moving towards that direction. 30, 67. In addition, the multireference diagnostic used in this study is by no means exhaustive, and investigation of multireference character in particular for a comparison of DLPNO-CCSD(T), AFQMC and other scalable methods is illustrative for the purposes of ascertaining the potential for evaluating challenging systems.

### v3

@@ Abstract
2.6 2.6. Basis Sets
We use the following basis sets: aug-cc-pVXZ-DK (X = T, Q) for atoms with atomic number less than or equal to that of oxygen, and aug-cc-pCVXZ-DK (X = T, Q) for fluorine and heavier, obtained using the Basis Set Exchange database.57, 58, 59, 60, 61, 62, 63 This choice is motivated by the documented improvement of basis set convergence when using core-valence or tight-d functions in the basis set for second-row elements (Na-Cl)64, 65, 66, 67 as well as fluorine68. While the split-valence aug-cc-pVXZ basis sets are not designed for core-valence correlation 69, we find that all-electron calculations using these same basis sets can reach a respectable (especially for CCSD(T)) albeit overall inferior accuracy (especailly for AFQMC) to the frozen core calculations supplemented with MP2 core corrections, as discussed briefly in Section 3, and in more in detail in SI Section 5.
We extrapolate all single point energies to the complete basis set (CBS) limit according to the method of Neese and Valeev70 for T/Q extrapolation for both Hartree-Fock and correlation energy, with $\alpha$ and $\beta$ matching the basis set used. We use the same coefficients $\alpha$ and $\beta$ for the core-valence sets aug-cc-pCVXZ-DK as the corresponding aug-cc-pVXZ-DK basis sets, where we use the aug-cc-pVXZ coefficients reported by Neese and Valeev. This CBS procedure is used by all the methods investigated, with the exception of the more expensive CCSD(T) and W-AFQMC where we use alternative schemes (see SI Section 3 and Section 6).
2.7 2.7. Frozen Core Corrections
@@
2.8 2.8. Relativistic Effects
Scalar relativistic effects are included through the DKH271, 72, 73 Hamiltonian for DLPNO-CCSD(T) and CCSD(T), and X2C74 for ph-AFQMC, as X2C is not implemented in ORCA and DKH2 is not implemented in PySCF. The MP2 core-valence corrections follow the same relativistic corrections for each method, respectively (MP2 is carried out in ORCA for correcting DLPNO-CCSD(T) and CCSD(T) corrections and in PySCF for ph-AFQMC corrections). The difference between DKH2 and X2C energies are shown to be small for HF and MP2 and mostly cancelled out by atomic energies (see SI Section 7). Spin-orbit corrections to atomic energies are applied using the values from Curtiss et al. 36.
2.9 2.9. Heat of Formation and Atomization
Atomization energies ($\sum E_{\text{atoms,{g}}}-E_{\text{molecule,{g}}}$) and heats of formation are calculated according to the method in Ref.36, 75, 76, including geometry optimization and thermochemical properties (ZPE, enthalpy, internal energy) using the DFT functional B3LYP 77, 78 and basis set 6-31G* 79, 80, 81, 82 using the Jaguar software package 83 with the maximum available grid point density. A few molecules required a higher-level geometry optimization (see SI Section 8). After optimization, DFT single-point energies were calculated with Jaguar. We note that the first step towards obtaining the heat of formation at 298 K is the atomization at 0 K, and the molecule temperature corrections (from DFT), atom temperature corrections (from experiment) and energy to change the atomic states from gas to standard state (from experiment, i.e. heat of formation of the single atom in the gaseous state) are added to achieve the heat of formation. Experimental values for the atomic heats of formation and temperature corrections are the same as that used by Curtiss et al.36, with the exception of the sulfur atom heat of formation where we use 66.18 kcal/mol from ATcT 42 instead of 65.66 kcal/mol.
@@
where $\Delta_{a}E(0\,\text{K})$ is the atomization energy at 0 K. With the addition of aformentioned temperature correction terms from DFT ($H^{298\text{ K}}-H^{0\text{ K}}$ for the molecule) and experiment ($H^{298\text{ K}}-H^{0\text{ K}}$ for the atom), we obtain $\Delta_{f}H(298\,\text{K})$. Although these corrections are approximate (even though we expect the errors to be small), they cancel out when obtaining the deviation from the converted W4 heat of formation reference value as the same corrections are applied to the calculated atomization energy of the molecule. Effectively, the deviation $D$ becomes
@@
2.10 2.10. Atomic Energies
The treatment of atoms is essential in achieving the targeted accuracy. An explicit near-exact treatment of core-valence correlation on par with valence correlation, as is done in W4-17, requires expensive core-valence corrections. In AFQMC, similar to other projection QMC approaches, the description of core-valence correlations requires onerous convergence of time-step errors. Furthermore, it has been shown that AFQMC atom energies can be difficult to calculate 33. Hence, we converge the atomic energies with a large number of determinants in W-AFQMC. A less expensive and simpler alternative is to fit the values of the atoms to the experimental data, which benefits from cancellation of errors. We use this approach for the other benchmark methods. For small molecules, a relatively inexpensive version of the first approach can be shown to work quite well, however, for larger molecules, any difference in molecule versus atom accuracy for a given method is compounded.
@@
where $\Delta_{f}H_{\text{expt},\text{molecule}}$ and $\Delta_{f}H_{\text{calc},\text{molecule}}$ are the experimental and calculated heats of formation for that molecule, and $N_{\text{atom}}$ is the number of the atom constituting the molecule. The energy of the atom, $E_{\text{atom}}$, is the fitting variable, and $c_{\text{molecule}}$ are the constant terms (for each molecule, i.e. independent of $E_{\text{atom}}$) such as the temperature terms, energy of the molecule, and heat of formation of the atoms, used when we minimize the left-hand side of the equation above. We have 189 such correlated equations for the 189 molecules in the G2 and G3 sets.
The initial guess for atom energies in the atomic energy fit is the ab initio atom energies obtained through each method respectively. It is worth noting that a small error (i.e. slight imperfection in cancellation of error between atom and molecule energy) in the ab initio atom energies is multiplied by the number of atoms in the molecule. While we do add spin-orbit corrections to atomic energies using the values from Curtiss et al. 36, this only applies to the initial guess and such atom-related corrections will be encompassed in the final fitted atom energy.
@@
3 3. Results and Discussion
3.1 3.1. Overall Performance of CCSD(T) and AFQMC-based Methods
We begin by evaluating the overall performance of the four methods discussed in Section 2: CCSD(T), DLPNO-CCSD(T), AFQMC I, and AFQMC II, across our chosen datasets. Figure 1 displays the root mean square deviation (RMSD) in the enthalpies of formation calculated using these methods, as detailed in Section 2. For the combined dataset (represented by the far-right bar in Figure  1), the RMSD values for all methods fall within 1 kcal/mol, and as we will see below, all four methods have a very small number of outliers with deviations from benchmark experiments or computations greater than 2 kcal/mol. We conclude that for main group chemistry, the localized (and hence scalable) versions of both coupled cluster and AFQMC achieve our target of reliably obtaining near-chemical accuracy for chemical transformations, sufficient for elucidating chemical reaction mechanisms in complex systems. Results along these lines have been presented previously for DLPNO-CCSD(T) (although not for as large and extensively curated a data set involving experimental and high-level theoretical references), but not for AFQMC. The present exercise establishes AFQMC as a robust alternative benchmark quantum chemical methodology, albeit at a higher computational cost than DLPNO-CCSD(T). For the present systems, we find the scaling exponent with system size to be similar between DLPNO-CCSD(T) and L-AFQMC, and the high prefactor is responsible for the cost of L-AFQMC. See SI Section 10 for examples of the computational costs of each method. We refer the reader to more detailed demonstrations of AFQMC and DLPNO-CCSD(T) scaling in these articles 24, 27, 10.
@@
Secondly, the most difficult data set for all methods is, unsurprisingly, but not guaranteed, the subset of 10 cases that we have identified as “multireference” (MR). We classify molecules as multireference based on a set of diagnostics developed by Karton et. al.40, 41, as discussed in more detail in Section 11 of the Supplementary Material. For the coupled-cluster based methods, only one molecule, ozone, stands out as displaying an error in excess of 2 kcal/mol. Despite formally being a single reference methodology, the treatment of electron correlation via CCSD(T) appears to be powerful enough to handle many wavefunctions with nontrivial multireference character. DLPNO-CCSD(T) here displays a noticeable (although not large) degradation from full CCSD(T). For AFQMC, an improvement is obtained in the treatment of MR molecules by upgrading the trial function in the AFQMC II approach. Details of results for each quantum chemical method for all of the MR cases can be found in the SI Section 11.
@@
We next analyze the outliers observed across the different data sets and methods, the number of which are summarized in Figure 2 below. A number of interesting points can be made regarding the outliers, which are enumerated along with results for the various computational methods in Table 2. Firstly, neither AFQMC I, AFQMC II nor the two coupled cluster based methods have any outliers for the W4-11 and W4-17 data sets, where the reference values are taken from ultrahigh level computation. Secondly, while CCSD(T) and DLPNO-CCSD(T) have a larger number of outliers than AFQMC II based on our (somewhat arbitrary) cutoff for the deviation from the reference of 2 kcal/mol, a perusal of the data in Table 2 shows that, with the exception of ozone (where we believe that the multireference character is great enough to cause significant errors in CCSD(T) and DLPNO-CCSD(T), which can then be reduced via a large trial function in AFQMC II), the computational results for the five remaining molecules are closer to each other than they are to the experimental reference data. A likely interpretation of the results is that the experiments have a residual error of a few kcal/mol (possibly as large as 3-4 kcal/mol for 3-butyn-2-one), and that in fact, the reliability of our scalable benchmark methods is higher than what is suggested in Figure 2. In contrast to a number of other cases that initially appeared to be outliers but were resolved by newer experiments, as discussed in Section 2.2, we were unable to find any relatively recent experimental data. Note that the value of having two distinct computational methods which can be compared, suggested in the Introduction, is already manifested in this analysis. Without the complementary AFQMC I and AFQMC II results, one might conclude that CCSD(T) has occasional outliers even for single-reference main group molecules and that higher-order treatments are required to achieve even the 2 kcal/mol accuracy threshold that we have set.
@@
3.2 3.2. Comparison of CCSD(T) and DLPNO-CCSD(T) Results
@@
3.3 3.3. Detailed Discussion of ph-AFQMC Methodology
In contrast to our coupled cluster based calculations, for which we utilize well established methods implemented by the Neese group in ORCA, the scalable AFQMC I protocol discussed above required significant novel methodology development. We therefore discuss AFQMC I development and implementation in detail in what follows. Specification of the AFQMC II protocol given AFQMC I as a starting point is straightforward, using the general procedure described in Section 2.4.
@@
Table 4: The top outliers from AFQMC 0 protocol. Deviations (($\Delta_{f}H_{\text{expt}}(298\,\text{K})-\Delta_{f}H_{\text{calc}}(298\,\text{K})$) for G2/G3 and --($\Delta_{a}E_{\text{W4}}(0\,\text{K})-\Delta_{a}E_{\text{calc}}(0\,\text{K})$) for W4, see Section 2.9) are listed in kcal/mol with statistical errors in parentheses. After the first CI is performed with an active space (AS) based on orbital maps to the atoms of the molecules (refer Table S4) that returns the ‘First CI AS’ listed, the second AS (shown here as ‘TZ final AS’ and ‘QZ final AS’, as the NOONs have a slight basis set dependency due to approximations such as the SHCI solver) is chosen from those orbitals from the first AS that have NOONs of between 0.01 and 1.99. The final number of determinants is set by the number of determinants required to get to 99.5% saturation of the CI coefficients.
@@
Under circumstances where the reference is unknown, typically having a converged energy with respect to, for example, determinants28 (See SI Section 3 for an example) gives confidence in the AFQMC benchmark value unless the CI expansion is qualitatively wrong. This process can become expensive, and based on our heuristic we have observed some guidelines for which type of trial and whether convergence is necessary for certain types of molecules. Observing the AFQMC 0 outliers, a few categories of molecules stand out: i) multireference molecules ii) small (< 4 atoms) molecules containing Li, F, or S atoms, iii) conjugated systems, iv) strained systems, and v) halogen oxoacids. The only exceptions to these are diethyl ketone (deviation 2.22 kcal/mol) and HNCO (deviation 2.63 kcal/mol). After an expanded valence space to two instead of one shell, and a 0.001 instead of 0.01 NOON threshold, the remaining real outliers as discussed above mostly fall into the multireference category. Therefore, for main group thermochemistry, for the above categories of outliers (except multireference), we recommend AFQMC 0 with these alternative thresholds. On the other hand, we still recommend that multireference molecules be converged with respect to the active space size and number of determinants.
@@
The G2 and G3 test sets and the W4 sets are chosen on account of the readily available and accurate reference values. Further critical investigation of scalable benchmark methods such as DLPNO-CCSD(T) and AFQMC for larger systems is valuable, with the difficulty of such investigation being the lack of accurate experimental references and the computational expense of generating benchmark-level calculations for large datasets of such systems. In particular, having the AFQMC method at disposal would be valuable in cases where one does not expect CCSD(T) to perform well, for example multireference systems or non-equilibrium geometries and bond breaking. Regardless, the current studies in the literature are moving towards that direction. 28, 70. The multireference diagnostic used in this study is by no means exhaustive, and investigation of multireference character for a comparison of DLPNO-CCSD(T), AFQMC and other scalable methods (for example, other L-CCSD(T) methods 84 and composite methods 85, 86) is illustrative for the purposes of ascertaining the potential for evaluating challenging systems.